Log I = Log 10^-16 + 9....Im having trouble solving for I. I understand Log 10 to any power is that number (in this case, -16). But I still am having trouble solving for I
1
Expert's answer
2012-08-07T08:59:10-0400
Log I = Log 10^-16 + 9
As you've mentioned absolutely correctly, Log 10^-16=-16*Log 10=-16
Therefore we have Log I = Log 10^-16 + 9=-16+9=-8
So due to definition of logarithm I=10^-8